Prove that the following two equations represent the simple harmonic motion-

X = A cosωt + B sinωt

X = A eiωt


1
Expert's answer
2021-05-08T14:39:06-0400

1)


"\frac{d^2x}{dt^2}=-\omega^2(A \cosωt + B \sinωt)=-\omega^2x\\\frac{d^2x}{dt^2}+\omega^2x=0"

2)


"\frac{d^2x}{dt^2}=-\omega^2(A e^{iωt}

)=-\omega^2x\\\frac{d^2x}{dt^2}+\omega^2x=0"